OpenStudy (barrelracer011):

HELP!!!!! Malia is using an online calculator to calculate the outputs f(n) for different inputs n. The ordered pairs below show Malia's inputs and the corresponding outputs displayed by the calculator. (1, 3), (2, 7), (3, 11), (4, 15) Which of the following functions best represents the rule that the calculator uses to display the outputs?

OpenStudy (barrelracer011):

f(n) = 5n - 1 f(n) = 5n + 1 f(n) = 4n + 1 f(n) = 4n - 1

OpenStudy (anonymous):

f(n)=4n-1